OpenStudy (anonymous):

Which of the following triangles are right triangles? Check all that apply A triangle with side lengths 5, 12, 13 A triangle with side lengths 6 inches, 8 inches, 12 inches A triangle with side lengths 8, 15, 17 A triangle with side lengths 9, 12, 15

OpenStudy (anonymous):

do you know?

OpenStudy (anonymous):

are you sure this is the answers below the question because i got different answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

because this is what i got A. A triangle with side lengths 4, 5, 6 B. A triangle with side lengths 6 inches, 8 inches, 10 inches but its not on yours

OpenStudy (anonymous):

oohh hmm i see. I honestly don't know.. but thanks..

OpenStudy (anonymous):

ok

OpenStudy (ankitshaw):

use pythagoras theorem to find out. if the sum of square of the smaller length are same as the square of the larger length then they form a right angle triangle. Like in option A the lengths given will form a right angle triangle 5^2+12^2=13^2 25+144 = 169 169 =169 ..... hope u understood!!

OpenStudy (ankitshaw):

so except second combination of lengths all the others form a right angle triangle